Two forces of 10 N each are acting on a point. One force acts east and the other 70 o north of east. Find the resultant displacement using the laws of sine and cosine.



1
Expert's answer
2021-09-06T11:35:08-0400

The angle between these two forces is "\theta = 70\degree". Accoring to the cosine law, the diagonal of the parallelogram formed by these two forces (the resultant force) is:


"F = \sqrt{10^2 + 10^2 + 2\cdot 10\cdot 10\cdot \cos70\degree} \approx 16N"

Answer. 16 N.